OpenStudy (anonymous):

What is the exact area of a circle whose diameter is 24 cm? A. 576 straight pi cm2 B. 144 straight pi cm2 C. 48 straight pi cm2 D. 24 straight pi cm2 @tkhunny @jackmullen55

OpenStudy (anonymous):

i got B. im i correct

OpenStudy (hysusonic):

do you know how to get diameter?

OpenStudy (hysusonic):

area* sorry

OpenStudy (anonymous):

yes

OpenStudy (hysusonic):

okay so half the diameter is the radius so then you plug in the numbers into pi*r^2 so 24/2=12 12^2=144pi

OpenStudy (anonymous):

pi*r^2 so 24/2=12 12^2=144pi. B. 144 straight pi cm2 A ≈ 452.39 cm²